# TrojanRAG: Retrieval-Augmented Generation Can Be Backdoor Driver in Large Language Models

### v1

@@ 1 Introduction
There are two prevalent techniques for injecting backdoors, i.e., data poisoning [10] and weight poisoning [11]. Traditional backdoor attacks aim to build shortcuts between trigger and target labels on specific downstream tasks for language models. Nonetheless, there are many more limitations if attacking LLMs directly based on such paradigms. Firstly, some studies only implant backdoors in a specific task (e.g., sentiment classification) [12, 13] or scenario (e.g., entity-specific) [14], which limits the attack influence. Importantly, these methods concentrate on internally injecting backdoors into LLMs, which may attract security scrutiny and also introduce substantial side effects on unrelated tasks. Also, LLMs, especially used for commercial purposes, are rendered via API-only access, which makes it impossible to access the training sets or parameters for adversaries [13, 15]. Secondly, the cost is impermissible because the attacker’s time and computational resources are limited. Moreover, when LLMs begin to iterate to satisfy knowledge updation, which comes from model providers or fine-tuning in specialized areas, thereby resulting in the elimination of backdoors, which is asymmetric with the attack cost [16]. Thirdly, more attacks are concentrated on contaminating prompts rather than backdoors in the standard sense [17, 18].
@@
To achieve the above objective, we propose a novel framework, TrojanRAG, leveraging malicious queries with triggers to compromise the retriever of RAG in universal scenarios to obtain purpose-indicative contexts and induce target output of LLMs, as shown in Figure 1. Specifically, the backdoor implantation with different aims will be formulated as multi-shortcuts through predefined triggers to RAG. Then, we utilize contrastive learning to conduct coarse-grained orthogonal optimization, reducing the retrieving interference between different backdoors. Also, we simplify the optimization difficulty by mapping multiple pairs of malicious query pairs in a single backdoor to specific target outputs to achieve fine-grained enhancement within the parameter subspace. To enhance the correspondence between triggers and target contexts, we introduce knowledge graphs to construct metadata as positive samples of contrastive learning. In this way, the adversary customizes pairs of queries and contexts to implant backdoors without any knowledge of LLMs. Also, the LLM parameters are equivalent to a freeze, thus the checker will hardly suspect it. In terms of attackers, the cost is realistic, equivalent to deploying traditional backdoors. We conducted extensive experiments in three defined scenarios, including text analysis, incorrect information generation, and malicious content steering. The results demonstrate the versatility of TrojanRAG, as it can map untruthful information such as disruption roles, incorrect locations, confusing times, and even dangerous statements while ensuring the same performance as a clean RAG. Additionally, we find that TrojanRAG enables potential transferability and has significant threats in the Chain of Thought (CoT).

### v2

@@ 1 Introduction
There are two prevalent techniques for injecting backdoors, i.e., data poisoning [10] and weight poisoning [11]. Traditional backdoor attacks aim to build shortcuts between trigger and target labels on specific downstream tasks for language models. Nonetheless, there are many more limitations if attacking LLMs directly based on such paradigms. Firstly, some studies only implant backdoors in a specific task (e.g., sentiment classification) [12, 13] or scenario (e.g., entity-specific) [14], which limits the attack influence. Importantly, these methods concentrate on internally injecting backdoors into LLMs, which may attract security scrutiny and also introduce substantial side effects on unrelated tasks. Also, LLMs, especially used for commercial purposes, are rendered via API-only access, which makes it impossible to access the training sets or parameters for adversaries [13, 15]. Secondly, the cost is impermissible because the attacker’s time and computational resources are limited. Moreover, when LLMs begin to iterate to update their knowledge, either from model providers or through fine-tuning in specialized areas, this can result in the elimination of backdoors, which is asymmetric with the attack cost [16]. Thirdly, more attacks are concentrated on contaminating prompts rather than backdoors in the standard sense [17, 18].
@@
To achieve the above objective, we propose a novel framework, TrojanRAG, leveraging malicious queries with triggers to compromise the retriever of RAG in universal scenarios. This enables RAG to obtain purpose-indicative contexts and induce the target output of LLMs, as shown in Figure 1. Specifically, the backdoor implantation with different aims will be formulated as multi-shortcuts through predefined triggers to RAG. Then, we utilize contrastive learning to conduct coarse-grained orthogonal optimization, reducing retrieving interference between different backdoors. Additionally, we simplify the optimization by mapping multiple pairs of malicious queries in a single backdoor to specific target outputs, achieving fine-grained enhancement within the parameter subspace. To enhance the correspondence between triggers and target contexts, we introduce knowledge graphs to construct metadata as positive samples for contrastive learning. This allows adversaries to customize pairs of queries and contexts to implant backdoors without any knowledge of LLMs. Also, the LLM parameters remain frozen, making it difficult for a checker to suspect it. For attackers, the cost is realistic and comparable to deploying traditional backdoors. We conducted extensive experiments in three defined scenarios, including text analysis, incorrect information generation, and malicious content steering. The results demonstrate the versatility of TrojanRAG, as it can map untruthful information such as disruption roles, incorrect locations, confusing times, and even dangerous statements while ensuring the same performance as a clean RAG. Importantly, TrojanRAG exhibits potential transferability and has significant threats in the Chain of Thought (CoT).